THE PURE ROTATIONAL SPECTRUM OF AlNC IN ITS X̃^{1}Σ^{+}(ν_{2} = 0, 1 and 2) STATES: A STRUCTURAL ANALYSIS

\underline{J. SCOTT ROBINSON}, Department of Chemistry and Biochemistry, Arizona State University,Tempe, AZ 85287-1604 USA; L.M. ZIURYS, Steward Observatory, University of Arizona 933 N Cherry Ave. Tucson, AZ 85721.

The pure rotational spectrum of aluminum isocyanide, AlNC, has been recorded at high resolution using millimeter-wave spectroscopy. The radical was produced in a flowing reactor/free space absorption cell by reacting aluminum vapor with a mixture of argon and cyanogen. The J = 10 through J = 31 rotational transitions in the fundamental mode and numerous features in the degenerate vibrationally excited bending mode (ν_{2}) were observed in the range 131-382 GHz. Spectroscopic parameters were derived from a least squares analysis of the transition frequencies. The derived constants are compared to theoretical calculations^{a}^{,}^{b}^{,}^{c}.

aBuyong Ma, Yukio Yamaguchi and Henry F. Schaefer III, Molec. Phys., 1995, 86(6), 1331.

bSimon Petrie, J. Phys. Chem., 1996, 100(28), 11581.

cColin Thomson, Int. J. Quantum Chem. Symp. 1976, 10, 85.